Write the Kuhn-Tucker conditions for the following problems and obtain the optimal solution:

Minimize                    Z= 2x1+3x2-x12- 2x22

Subject to      x1+3x2 ≤ 6,

 5x1+2x2 ≤ 10,

               x1, x2≥ 0


1
Expert's answer
2022-01-14T05:18:02-0500

1) "g_{i}(x^*)-b_{i}" is feasible for "i=1,2....m"


Where "g_{i}(x^*)" is optional value


Writing the equation in residue form


"f)x)=2x_{1}+3x_{2}-x_{1}^2-2x_{2}^{2}"


"g_1(x)=x_1+3x_2-6"


"g_2(x)=5x_1+2x_2-10"



Applying the Kuhn-Tucker conditions (Optimality Condition)


2)


"\begin{pmatrix}
 \frac { \partial\>f}{\partial\>x_1} \\\\
 \frac{\partial\>f}{\partial\>x_2} 
\end{pmatrix}" "+\lambda1" "\begin{pmatrix}
 \frac{\partial\>g_1}{\partial\>x_1} \\\\
 \frac{\partial\>g_1}{\partial\>x_2} 
\end{pmatrix}" "+\lambda2" "\begin{pmatrix}
 \frac{\partial\>g_2}{\partial\>x_1} \\\\
 \frac{\partial\>g_2}{\partial\>x_2} 
\end{pmatrix}=0"



"2-2x_1+\lambda_1+5\lambda_2=0......(1)"

"3-4x_2+3\lambda_1+2\lambda_2=0.....(2)"



Feasibility condition


"x_1+3x_2\le6" "......(3)"

"5x_1+2x_2\le10\>\>......(4)"


Complimentary slackness property


"\lambda_1(x_1+3x_2-6)=0" "....(5)"

"\lambda_2(5x_1+2x_2-10)=0\>.....(6)"



Non-negativity constraints


"\lambda_1\ge0,\>\>\lambda_2\ge0\>\>.....(7)"



Case (i)


"\lambda_1=0,\lambda_2=0"

From (1) and(2), "2-2x_1=0\implies\>x_1=1"

"3-4x_2=0\implies\>x_2=\frac{3}{4}"


Point "(1,\frac{3}{4})" is a KKT point


Case (ii)


"\lambda_1=0,\lambda_2\ne0"

Using (1),(2) and (6)

"2-2x_1+5\lambda_2=0"

"3-4x_2+2\lambda_2=0"

"5x_1+2x_2-10=0"


"x_1=\frac{2+5\lambda_2}{2},\>\>x_2=\frac{3+2\lambda_2}{4}"


"5(\frac{2+5\lambda_2}{2})+2(\frac{3+2\lambda_2}{4})=0"


"54\lambda_2=14"

"\lambda_2=\frac{7}{27}"


"x_1=\frac{2+5(\frac{7}{27})}{2}=\frac{89}{54}"


"x_2=\frac{3+2(\frac{7}{27})}{4}=\frac{95}{108}"


This could be a KKT point


Case(iii)


"\lambda_1\ne0,\lambda_2=0"

"2-2x_1+\lambda_1=0"

"3-4x_2+3\lambda_1=0"

"x_1+3x_2-6=0"


"x_1=\frac{2+\lambda_1}{2}" "x_2=\frac{3+3\lambda_1}{4}"


"\frac{2+\lambda_1}{2}+\frac{3(3+3\lambda_1)}{4}=6"

"\implies\>\lambda_1=1"


Substituting in "x_1" and "x_2"

"x_1=\frac{3}{2},\>\>x_2=\frac{3}{2}"


But "5x_1+2x_2=10.5"

This violate ......(4)


"\therefore(\frac{3}{2},\frac{3}{2})" is not a KKT point.



Case (iv)


"\lambda_1\ne0,\>\lambda_2\ne0"


"2-2x_1+\lambda_1+5\lambda_2=0\>......(i)"

"3-4x_2+3\lambda_1+2\lambda_2=0\>....(ii)"

"x_1+3x_2=6\>.....(iii)"

"5x_1+2x_2=10\>.....(iv)"



Solving (iii) and (iv)


"5x_1+15x_2=30"

"5x_1+2x_2=10"

"\implies13x_2=20"

"x_2=\frac{20}{13}"

"x_1=\frac{18}{13}"


Substituting "x_1" and "x_2" in (i) and (ii)


"\lambda_1+5\lambda_2=\frac{10}{13}"

"3\lambda_1+2\lambda_2=\frac{41}{13}"


Solving

"3\lambda_1+15\lambda_2=\frac{30}{13}"

"3\lambda_1+2\lambda_2=\frac{41}{13}"

"\lambda_2=\frac{-11}{169},\>\lambda_1=\frac{185}{169}"


"\lambda_2<0"

This violates (7)


"z=2x_1+3x_2-x_1^2-2x_2^2"


Substituting point "(1,\frac{3}{4})"


"z=2(1)+3(\frac{3}{4})-(1)^2-2(\frac{3}{4})^2"

"=2.125"



Substituting point "(\frac{89}{54},\frac{95}{108})"


"z=2(\frac{89}{54})+3(\frac{95}{108}-(\frac{89}{54})^2-2(\frac{95}{108})^2"


"=1.671296"


Minimum "z=1.671296"

at point "(\frac{89}{54},\frac{95}{108})"